# Point3R: Streaming 3D Reconstruction with Explicit Spatial Pointer Memory

### v1

@@ 5 Conclusion and Discussions
In this paper, we have presented an online streaming 3D reconstruction framework, Point3R, with a spatial pointer memory. When processing streaming inputs, our method maintains a growing spatial pointer memory in which each pointer is assigned a specific 3D position and aggregates scene information nearby with a changing spatial feature. Equipped with a 3D hierarchical position embedding and a simple yet effective memory fusion mechanism, our method imposes minimal constraints on the input, handling both static and dynamic scenes as well as ordered or unordered image collections. With a low training cost, our method achieves competitive or state-of-the-art performance on various 3D/4D tasks, which verifies the effectiveness of our method.
@@
Broader Impacts. Our method facilitates scalable and efficient dense streaming 3D scene reconstruction, benefiting a wide range of applications. The explicit and interpretable design of our pointer memory makes our method more transparent and adaptable to different real-world scenarios.
@@ A.1 Memory Fusion Mechanism
For $n$, $m$-th query and key $\mathbf{q}_{n},\mathbf{k}_{m}\in\mathbb{R}^{1\times d_{head}}$, RoPE converts them to complex vector $\bar{\mathbf{q}}_{n},\bar{\mathbf{k}}_{m}\in\mathbb{R}^{1\times(d_{head}/2)}$ by considering ($2t$)-th dim as real part and ($12t+$ 1)-th dim as imaginary part. We follow this so $\theta_{t}$ in Eq. 14 in the main text is
@@
Table 6: Training Datasets.
Appendix B More Implementation and Training Details
Table 6 provides more details about the datasets we use to train our model. To be specific, we use ARKitScenes, BlendedMVS, CO3Dv2, HyperSim, MegaDepth, ScanNet, ScanNet++, WildRGBD, VirtualKITTI2 and Waymo datasets in the first stage which takes almost 2 days. We add other datasets and increase the input resolution in the second stage which takes almost 2 days. The final stage (longer input sequences) takes about 3 days. Besides, we do not use the memory fusion mechanism in the first stage to improve the stability of training. We dynamically adjust the usage ratio of each dataset at different stages during our training to improve the overall model performance.
Appendix C More Visualizations

### v2

@@ 5 Conclusion and Discussions
In this paper, we have presented an online streaming 3D reconstruction framework, Point3R, with a spatial pointer memory. When processing streaming inputs, our method maintains a growing spatial pointer memory in which each pointer is assigned a specific 3D position and aggregates scene information nearby with a changing spatial feature. Equipped with a 3D hierarchical position embedding and a simple yet effective memory fusion mechanism, our method can handle both static and dynamic scenes as well as ordered or unordered image collections. With a low training cost, our method achieves competitive or state-of-the-art performance on various 3D/4D tasks.
@@
Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB280693, in part by the National Natural Science Foundation of China under Grant 62125603, Grant 62336004, Grant 62576188, and Grant 62321005, in part by the Beijing Natural Science Foundation under Grant L247009, and in part by the Beijing National Research Center for Information Science and Technology.
@@ A.1 Memory Fusion Mechanism
For $n$, $m$-th query and key $\mathbf{q}_{n},\mathbf{k}_{m}\in\mathbb{R}^{1\times d_{head}}$, RoPE converts them to complex vector $\bar{\mathbf{q}}_{n},\bar{\mathbf{k}}_{m}\in\mathbb{R}^{1\times(d_{head}/2)}$ by considering ($2t$)-th dim as real part and ($2t+1$)-th dim as imaginary part. We follow this so $\theta_{t}$ in Eq. 14 in the main text is
@@
A.3 Pose Retrieval Mechanism
For each input sequence, we introduce a learnable token that is initialized within the model and serves as the pose token of the first frame. This token acts as the latent representation of the camera pose at the beginning of the sequence and provides a reference for subsequent frames. In addition to this initialization, we define a compact set of learnable tokens that function as a lightweight pose memory module [63]. This memory is specifically designed for pose retrieval, enabling the model to accumulate and reuse pose-related information across frames in a sequence without incurring excessive computational cost.
After processing each frame, the pose token that has interacted with the current image features is used to update this memory set, thereby maintaining a condensed yet informative representation of previously observed poses. When the next frame arrives, its corresponding pose token is initialized not randomly, but by retrieving an initial value from this pose memory. This retrieval process leverages the image tokens of the current frame to query the memory and select the most relevant pose context. In this way, the proposed memory mechanism allows the model to efficiently capture temporal dependencies between frames while avoiding redundancy and instability.
Table 9: Training Datasets.
@@ A.3 Pose Retrieval Mechanism
Table 10: Robustness to poor initialization frames.
Normal Initialization Poor Initialization
[…]
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
CUT3R [63] 0.166 0.040 0.081 0.016 0.980 0.976 0.290 0.154 0.156 0.051 0.733 0.942
Ours 0.096 0.037 0.058 0.014 0.977 0.977 0.120 0.070 0.068 0.019 0.853 0.972
Appendix B More Training Details
Table 9 provides detailed information about the datasets used to train our model. To ensure both diversity and robustness, we adopt a multi-stage training strategy that progressively enhances the model’s generalization capability across a wide range of real-world and synthetic environments.
In the first stage, we train our model on a collection of large-scale datasets, including ARKitScenes, BlendedMVS, CO3Dv2, HyperSim, MegaDepth, ScanNet, ScanNet++, WildRGBD, VirtualKITTI2, and Waymo. Such diversity allows the model to learn robust geometric priors and generalizable visual representations that are independent of specific dataset characteristics.
In the second and third stage, we introduce additional datasets and increase the input resolution. This progressive training paradigm enables the model to gradually adapt from coarse-level scene reasoning to high-resolution, detail-oriented prediction, effectively improving its performance in both accuracy and stability.
To further stabilize training in the early phase, we disable the memory fusion mechanism during the first stage. This prevents unstable memory updates that may arise when both pose and feature representations are still under training. Once the model achieves stable convergence, the memory fusion mechanism is introduced to facilitate temporal consistency learning across frames and improve the model’s efficiency.
Moreover, we employ a dynamic dataset sampling strategy, where the sampling ratio of each dataset is adaptively adjusted throughout training. This adaptive reweighting ensures that the model receives balanced supervision from different datasets, mitigates potential domain bias, and leads to improved generalization across diverse scene distributions.
Appendix C More Experiments
Robustness to Poor Initialization Frames. The use of the initial frame as the reference leading to possible sensitivity to poor initialization frames is a common issue faced by the “-3R” series of work. We therefore evaluate the robustness of our method on such scenarios. We select Scene Kitchen in the NRGBD dataset and obtain 37 input images by sampling at the interval of 40 frames. To simulate a poor initialization, we set the 23rd sample (corresponding to frame 920 in the original dataset, which mainly contains large areas of the low-texture wall and floor with no other objects) as the first frame and randomly shuffled the order of the remaining frames. The comparison of results between our method and CUT3R under these two input settings is shown in Table 10.
We observe that our method is more robust to the effect of the initialization. Although the performance slightly degrades when the initial frame has significantly degraded content, the drop is acceptable considering the difficulty of such a scenario. We think this is because our explicit pointer memory is aligned with 3D structures of the current scene, which can minimize the impact of initialization as much as possible.
Appendix D More Visualizations
@@
Appendix E Broader Impacts
Our method facilitates scalable and efficient dense streaming 3D scene reconstruction, benefiting a wide range of applications. The explicit and interpretable design of our pointer memory makes our method more transparent and adaptable to different real-world scenarios.